1. Convert the polar coordinates (-8, 2π/3) into rectangular coordinates.
  2. Convert the rectangular coordinates (3, -3) into polar coordinates with r > 0 and 0 ≤ θ < 2π.
  3. Convert the rectangular equation x2 + y2 = 100 into a polar equation that expresses r in terms of θ.
  4. Convert the polar equation 4r cos θ + r sin θ = 8 into a rectangular equation that expresses y in terms of x.
1
Expert's answer
2021-12-02T07:20:48-0500

Let "(r, \theta)" - polar coordinates and "(x, y)" - rectangular coordinates.


1) "r_{1} = - 8"

"\theta_{1} = \dfrac{2\pi}{3}"

"x_{1} = r_{1}cos(\theta_{1}) = (-8) * cos \left(\dfrac{2\pi}{3}\right) = (-8) * \left(-\dfrac{1}{2}\right) = 4"

"y_{1} = r_{1}sin(\theta_{1}) = (-8) * sin \left(\dfrac{2\pi}{3}\right) = (-8) * \left(\dfrac{\sqrt{3}}{2}\right) = -4\sqrt{3}"

2) "x_{2} = 3"

"y_{2} = - 3"

"r_{2} > 0"

"0 \le \theta_{2} < 2\pi"

"r_{2}^{2} = x_{2}^{2} + y_{2}^{2} = 3^{2} + (-3)^{2} = 18 \rArr r_{2} = 3\sqrt{2}"

"tan(\theta_{2}) = \dfrac{y_{2}}{x_{2}} = -1"

"x_{2} > 0, y_{2} < 0 \rArr \theta_{2} = \dfrac{7\pi}{4}"


3)"x^{2} + y^{2} = 100"

"r^{2} = 100 \rArr r = 10"


4) "4 r cos\theta + rsin \theta = 8"

"4x + y = 8"

"y = - 4x + 8"